OpenStudy (anonymous):

a sheet of standard-size copy paper measures 8.5 in. by 11 in. if a ream (500 sheets) of this paper has a volume of 187 in. how thick is the ream

OpenStudy (anonymous):

volume = Length*width*height V= 8.5*11*h 187=8.5*11*h h=187/ 8.5*11

OpenStudy (anonymous):

h=2

OpenStudy (anonymous):

thats the thickness ;)

OpenStudy (anonymous):

Thank you!